If Tan (theta)=7/24 and (theta) is acute, find the exact values of sec (theta), cosec (theta) and cot (theta)
Answer by lwsshak3(11628)

If Tan (theta)=7/24 and (theta) is acute, find the exact values of sec (theta), cosec (theta) and cot (theta)
***
Reference right triangle is in quadrant I with opposite side=7, adjacent side=24
hypotenuse=√((7^2)+(24^2))=√(49+576)=√625=25
..
sec(theta)=25/24
csc(theta)=25/7
cot(theta)=24/7